At 4.75 percent interest, how long does it take to quadruple your money?
At 4.75 percent interest, how long does it take to double your money?

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

(1)these questions can't be answered because you don't have an amount to add to the interest,it could be any amount

Answer by Alan3354(69443)

At 4.75 percent interest, how long does it take to quadruple your money?
At 4.75 percent interest, how long does it take to double your money?
-----------------
You didn't give the compounding period, if any.
Compounded annualy:
2*Amt = Amt*1.0475^t (t in years)
1.0475^t = 2
t*log(1.0475) = log(2)
t = log(2)/log(1.0475)
t =~ 14.936 years to double
-------------
Do the others the same way.